OpenStudy (anonymous):

What is the slope and the y-intercept for the line represented by this equation? A. B. C. D.

OpenStudy (anonymous):

what equation

OpenStudy (haseeb96):

@asdfsadfsadf where is the equation ???

OpenStudy (anonymous):

y=3/8x-13

OpenStudy (anonymous):

3/8 is the slope because the numbers before x always equal slope

OpenStudy (haseeb96):

y=mx + b this is slope intercept formula now in this above equation m is the slope and b is the y intercept in order to find slope and y intercept so just compare the Question equation with this equation y =mx +b

OpenStudy (haseeb96):